# DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition of Human Activities in Smart Homes

### v1

@@ 6.2. Results
[…]
Table 4. A side-by-side comparison of original CASAS labels and the more granular sub-activities identified by DISCOVER in Milan, Aruba, and Cairo. CASAS Labels are mapped to DISCOVER labels. Dashes indicate that this activity cannot be identified in that household (e.g., Aruba and Cairo have no bathroom sensors). These results showcase how DISCOVER is able to discover much more granular sub-activities for all three CASAS datasets.
We now want to explore these manually annotated sub-activity labels in greater detail. Table 4 illustrates how each discovered sub-activity label maps onto the broader CASAS categories (e.g., Cook, Relax, Sleep), along with the percentage of time-window sequences $W_{i}$ that each label occupies in Milan, Aruba, and Cairo. Dashes indicate activities that cannot occur in a given household (e.g., no guest bathroom sensors in Aruba or Cairo).
Notably, the total proportion of each broad CASAS label usually aligns reasonably well with our sub-activity categories. In Milan, for example, Sleep comprises around 7% of the CASAS data and 8% of our sub-activity labels (movement in bedroom, movement near bed). Similarly, CASAS marks 41% of the Milan data as Relax, compared to 36% for our corresponding sub-activities. More importantly, our approach results in a much finer breakdown of these broad labels, such as distinguishing sitting in living room armchair vs. sitting in office armchair, or movement near the medicine cabinet vs. movement near the stove. In Aruba, a single Relax label can be further divided into sitting on the couch vs. motion in a TV chair. These finer distinctions highlight one of the key advantages of DISCOVER: while it can reasonably well replicate higher-level annotations from existing datasets, it can also uncover more granular activities that may carry significant behavioral or clinical relevance.
@@ 7. Discussion
In this section, we discuss how DISCOVER can be used in real-world deployable applications (Section 7.1). We also outline the limitations of our approach and propose directions for future research (Section 7.2).
7.1. Potential Use Cases
In this section, we illustrate the real-world potential of our approach by demonstrating its ability to automatically annotate ambient sensor datasets and re-annotate existing data at finer granularity levels. Finally, we provide some precise implementation details, including a minute-by-minute reconstruction of daily activity patterns.
7.1.1. Annotation of Massive Unlabeled Ambient Sensor Datasets:
@@
Smart homes have been studied extensively across academia, industry and medical fields. While fully annotated datasets such as CASAS (Cook et al., 2009), Marble (Arrotta et al., 2022), and Orange4Home (Cumin et al., 2017) are widely used in AI research, they are far from the norm. Large-scale medical initiatives such as CART (Beattie et al., 2020) (272 homes over 3 years) and TIHM (Palermo et al., 2023) (56 homes over 50 days) produce orders of magnitude more data, but remain less commonly used due to their lack of activity labels resulting from prohibitively high costs of manual annotation. With DISCOVER, it becomes feasible to automatically annotate such massive datasets. Users can choose their desired level of granularity and exercise control over the annotation process via an interactive visualization tool that displays sensor activation sequences for each cluster. While such labels may not constitute true ground truth, which is inaccessible without camera data or similar information, our analysis has shown that processing with DISCOVER results in semantically meaningful activity clusters that can be labeled with high degree of agreement, thus providing insight into common everyday activities. We believe that DISCOVER can expand access to previously unlabeled data, creating opportunities for advancement in the areas such as in-home health monitoring (Morita et al., 2023), early detection of Alzheimer’s (Alberdi et al., 2018), and intelligent support for individuals with dementia (Demir et al., 2017).
7.1.2. Re-annotation of Existing Datasets with Finer Granularity:
@@
As we have seen in case of CASAS (Cook et al., 2012), existing datasets in the HAR domain often provide only coarse labels that capture broad activities, obscuring the rich variability of human behavior. DISCOVER addresses this shortcoming by enabling the discovery of sub-activities and variants within the data. For instance, in sleep monitoring, rather than merely recording an eight-hour sleep duration, our method could allow to differentiate between distinct sub-activities such as tossing and turning, or identifying periods of nocturnal bathroom visits. Similarly, in meal preparation, instead of labeling the entire process as Cooking, our approach can reveal detailed steps such as time spent at the stove, visits to the pantry, and travel to/from the dining room. By re-annotating existing datasets with these finer-grained labels, DISCOVER provides deeper insights into behavioral patterns, which can be particularly valuable in clinically relevant applications and personalized intervention strategies.
7.1.3. Temporal Sequence Analysis
Figure 6. A minute by minute view of atomic activities in the Milan household for two hours. Each colored segment represents a specific atomic activity (e.g., movement near stove, entering guest bathroom), allowing for a fine-grained understanding of daily routines and transitions. This higher-resolution approach goes beyond broad categories like cooking or bathing, highlighting detailed behaviors and potential health or lifestyle insights.
Building on the above, annotation of HAR data at greater granularity facilitates more detailed temporal analysis of human behavioral data. Specifically, our cluster-based approach facilitates the reconstruction of a continuous temporal record of daily activities by assigning each time window to a specific cluster. As illustrated in Fig. 6, a minute-by-minute snapshot of a Milan household on a given date displays various sub-activity labels (e.g., movement near stove, entering guest bathroom). The color scale represents the model’s confidence in cluster assignments, transitioning from low confidence (green) to high (red). This figure reveals coherent temporal sequences, such as the resident remaining in the living room armchair from 9:00 to 9:15 AM before transitioning to kitchen activities and entering the guest bathroom around 9:25 AM. Lower-confidence activities, often occurring during transitions between rooms, can be filtered out to produce a more stable activity timeline.
By delineating detailed sequences of sub-activities, our method enables domain experts and researchers to gain deeper insights into residents’ daily routines. This granularity allows for the identification of significant behavioral changes, such as increased bathroom visits indicating potential health issues or prolonged armchair sitting reflecting shifts in mobility or energy levels, factors that are critical for in-home health monitoring (Morita et al., 2023). Additionally, the continuous temporal component supports real-time interventions, adaptive healthcare monitoring, and personalized smart home automations.

### v3

@@ 5.2. Evaluation of Consensus and Cluster Homogeneity
Milan Aruba Cairo CASAS Label DISCOVER Label CASAS DISCOVER CASAS DISCOVER CASAS DISCOVER Cook TOTAL 31% 19% 39% 24% [0.5pt/2pt] Movement all over kitchen 4% 24% Movement near Fridge Movement near Stove 2% Movement near Medicine Cabinet 13% Eat TOTAL 2% 6% 4% 6% 72% 67% [0.5pt/2pt] Movement all over Dining Room Movement through Kitchen + Dining 6% 6% 32% Movement through Living + Dining 35% Relax TOTAL 41% 36% 49% 37% 0% 7% [0.5pt/2pt] Motion all over Living room 0% Motion in TV chair and area 6% Sitting on couch/armchair 19% 29% 7% Sitting in office armchair 15% Movement through Kitchen + Living 2% 2% Work TOTAL 3% 4% 4% 7% 12% 0% [0.5pt/2pt] Movement near office computer/desk 4% 7% Sleep TOTAL 7% 8% 3% 3% 12% 7% [0.5pt/2pt] Motion in all over Master Bedroom 8% 0.1% In Master Bedroom, Movement near bed 3% 7% Bathing TOTAL 14% 19% [0.5pt/2pt] Leaving the Guest Bathroom Movement inside the Guest Bathroom Entering Master Bathroom 1% Leaving Master Bathroom Movement in Master bathroom 5% Leave_Home Motion near entrance doors 2% 6% 0.3% 6% 1% 5% Motion through house 0% 4% 0% 14% 3% 13% TOTAL 100% 100% 100% 100% 100% 100%
Table 3. Comparison of broad CASAS labels and granular DISCOVER patterns across three datasets. Percentages show the distribution of labels; dashes (-) indicate activities unsupported by a home’s specific sensor topology (e.g., no bathroom sensors in Aruba and Cairo). The results highlight DISCOVER’s ability to identify finer behavioral distinctions while avoiding the semantic rigidity of fixed supervised categories.
5.3. Comparative Analysis: DISCOVER PDLs vs. CASAS Annotations
In this section, we explore the relationship between the identified patterns of daily living (PDL) and the original CASAS annotations. Table 3 illustrates how each discovered PDL label maps onto broader CASAS categories (e.g., Cook, Relax, Sleep), along with the percentage of time-window sequences $W_{i}$ each label occupies. While the overall distribution of activities is consistent across both protocols, DISCOVER offers two distinct advantages: i) finer granularity; and ii) contextual flexibility. First, DISCOVER provides a highly specific breakdown of broad categories. For example, in Milan, we distinguish between sitting in a living room armchair versus an office armchair, and in Aruba, a single Relax label is further deconstructed into sitting on the couch versus motion in a TV chair. Second, and more importantly, DISCOVER avoids the semantic rigidity inherent in supervised or transfer learning approaches. In those traditional approaches, a physical behavior (e.g., “sitting in an office armchair”) is often tied to a fixed label (e.g., Work) based on source-home assumptions. In contrast, DISCOVER remains agnostic to such priors, allowing the same physical location to map to different behaviors based on actual usage: in Milan, the office armchair is associated with Relaxation, because that’s the best place in house to watch the TV from, while in Aruba, it is associated with Work. By solely relying on the underlying data patterns without imposing external definitions, DISCOVER enables more robust longitudinal trend tracking. This allows researchers to first establish a household’s unique habits and subsequently detect meaningful behavioral shifts over time without the bias of predefined activity mappings.
5.4. Visualizing PDLs using t-SNE
Figure 4. tSNE projection of SCAN embeddings from the Milan household; each point represents a sensor window embedding colored by its original CASAS label. Insets display the deployment environment layout overlaid with a heatmap of sensor activations. Insets (a) and (b) highlight two distinct clusters within the CASAS Cook label—cluster 16 showing movement between kitchen and dining areas, and cluster 5 capturing activity near the medicine cabinet. Insets (c) and (d) show clusters from the Relax label, corresponding to sitting in the TV room armchair and sitting in the living room armchair. Data associated with the Other label (gray) is dispersed across clusters, underscoring its heterogeneous nature. This figure showcases the capability of DISCOVER to uncover more granular and nuanced behavioral patterns than the original CASAS labels.
In this section, we perform a visual inspection of Milan clusters in order to gain more insights into the generated clusters. Fig. 4 shows a tSNE (Maaten and Hinton, 2008) projection of the SCAN embeddings from the Milan dataset; each point represents a 768-dimensional sequence embedding colored by its original CASAS annotation. While some activities (e.g., Work) form single clusters, data representing other activity labels (e.g., Cook, Relax) fall into multiple clusters. Each point in the scatterplot is annotated with a probability score reflecting SCAN’s confidence in assigning that sequence to its respective cluster, offering interpretability and a level of control that simpler clustering methods (e.g., k-means) lack.
Fig. 4 insets take a closer look at two Cook and two Relax clusters. Each inset includes the layout of the deployment environment overlaid with a heatmap representing sensor activation patterns for the data sequences associated with that cluster. Insets (a) and (b) are associated with the Cook CASAS label (orange). Although both sets of data are labeled as Cook within CASAS, we observe clear differences between the sensor patterns within each cluster. Cluster 16 (inset (a)) captures the person leaving the kitchen and moving to the dining room, with motion sensors from the dining room and part of the dining room registering the person’s presence. By comparison, cluster 5 (inset (b)) is tightly focused on activity around the medicine cabinet in the kitchen. Insets (c) and (d) are associated with the Relax CASAS label (burgundy), where cluster 3 (inset (c)) captures a static activity in the TV room armchair while cluster 17 (inset (d)) corresponds to sitting in a living room armchair instead. This visualization illustrates a key advantage of DISCOVER: its ability to reveal household-specific, fine-grained patterns of daily living, such as sitting in different armchairs or engaging in distinct kitchen routines, that are difficult to capture using predefined activity labels. In contrast, fully supervised or transfer-based HAR approaches typically collapse such behaviors into coarse categories (e.g., Relax or Cook), limiting their ability to represent individualized behavioral structure.
6. Results: Quantitative Evaluation and Sensitivity Analysis
In this section, we ignore the DISCOVER human annotator labels, and evaluate only the quality of the clusters generated in Steps 0-2 of our approach. Examining the clusters helps provide further insights into the degree to which they match previous hand-annotated labels from CASAS, and how they are able to offer an even finer level of detail than the original annotations did. To validate cluster performance, we use CASAS as a ground truth oracle to temporarily assign a label to each cluster through majority voting. We then perform a direct comparison with fully supervised approaches, and investigate the impact of varying cluster count, examining the trade-off between granularity, label alignment and annotation costs.
6.1. Performance Alignment with Supervised Baselines
Model Method Type Milan Aruba Cairo
DeepCASAS Supervised (pre-seg) 0.79 0.93 0.82
TDOST Supervised (pre-seg) 0.83 0.92 0.82
Tree-CNN Supervised (no pre-seg) 0.65 0.77 0.76
DISCOVER Self-Supervised (no pre-seg) 0.82 0.90 0.85
DISCOVER vs Best -0.01 -0.03 +0.03
Table 4. Benchmarking DISCOVER against SoTA fully-supervised baselines using weighted F1 score. Bold values indicate the best performance for each dataset. The final row reports the absolute gap between the best supervised method and DISCOVER. DISCOVER demonstrates near SoTA results (and surpassing Cairo) on non pre-segmented data.
To evaluate the quality of the discovered clusters, we assess their alignment with established CASAS labels. Since DISCOVER identifies Patterns of Daily Living (PDL) without prior knowledge of these labels, we utilize a majority-vote mapping to assign a semantic category to each cluster based on its constituent samples. This approach allows us to directly compare our self-supervised framework against state-of-the-art fully supervised pipelines: DeepCASAS (Liciotti et al., 2020), TDOST (Thukral et al., 2024), and TreeCNN (Cao et al., 2020). Following previous papers (Thukral et al., 2024; Demirel et al., 2018; Bouchabou et al., 2021b, c; You et al., 2025a), we utilize the weighted F1 metric to evaluate classification performance, ensuring our results are directly comparable to established benchmarks. DeepCASAS is a fully supervised deep learning pipeline that leverages LSTM architecture to capture temporal dependencies in the data. TDOST is transfer learning pipeline that leverages full supervision on the source dataset. TDOST converts sensor triggers into natural language descriptions using LLMs, which provides the model with contextual information and enables the transfer of a single model across multiple households. Both DeepCASAS and TDOST rely on pre-segmented activity windows (for both source and target datasets in case of TDOST). TreeCNN is the last supervised baseline that leverages tree-structured convolutional neural network to. Similar to DISCOVER, TreeCNN does not rely on pre-segmentation and therefore serves as a suitable baseline. While surpassing supervised models is not the primary objective of a discovery-focused framework, a close performance alignment signifies a reasonable alignment between our PDLs and CASAS labels.
For a rigorous comparison, we re-trained DeepCASAS and TDOST on continuous sliding windows ($W_{i}$) to match the realistic, non-pre-segmented operating conditions of DISCOVER. As shown in Table 4, DISCOVER demonstrates remarkable performance alignment with these baselines. The performance delta, reported in the final row as “DISCOVER vs Best,” reveals that the gap is minimal across all households. In the Milan and Aruba datasets, DISCOVER remains within 3 percentage points (pp) of the best supervised results. Notably, in the Cairo dataset, DISCOVER outperforms all supervised baselines, achieving an F1 score of 0.85 compared to the 0.82 achieved by DeepCASAS and TDOST.
These results underscore the high fidelity of our self-supervised embeddings. Without the benefit of training labels, DISCOVER organizes raw sensor triggers into clusters that are not only semantically coherent but also match or exceed the classification accuracy of models requiring exhaustive manual annotation. These findings suggest that the latent structures discovered by DISCOVER are fundamentally aligned with the canonical activities traditionally defined by human annotators, while offering the added benefit of discovering the nuanced patterns discussed in Section 2.1.
6.2. The Impact of Representation-Aware Clustering
Method Milan Aruba Cairo
K-Means 0.65 0.66 0.66
SCAN 0.79 0.93 0.82
Table 5. Comparison of K-Means and SCAN clustering performance using F1 score. Bold values indicate the best performance for each dataset. SCAN consistently outperforms K-Means across all datasets, demonstrating the importance of representation-aware fine-tuning.
In prior sections, we motivated SCAN as our primary clustering model because it fine-tunes the encoder to enforce local embedding consistency among nearest neighbors. In this section, we want to compare SCAN clusters to a simpler geometric method such as k-Means applied to learned BERT embeddings.
Table 5 shows how SCAN F1 scores reported in Section 6.1 compare to k-Means clusters. Across all three CASAS datasets, k-Means performs substantially worse than SCAN on Weighted F1 scores. This behavior is expected: ambient smart-home data is sparse and highly heterogeneous Pure geometric clustering fails to recover rare behaviors and often condenses multiple unrelated patterns of daily living into the same cluster. As a result, K-Means is unable to form coherent and interpretable patterns of daily living in the way SCAN does. By contrast, SCAN’s fine-tuning phase encourages clusters whose neighbors share similar context, leading to significantly more stable and semantically meaningful clusters. Representation-aware objectives such as SCAN’s consistency loss are essential for discovering recurring routines in smart-home environments.
Figure 5. F1 scores for varying numbers of SCAN clusters ($k$) on Milan, Aruba, and Cairo datasets. The chart shows average F1 scores with bootstrapped 95% confidence intervals . Increasing $k$ up until 20-40 clusters improves alignment with CASAS labels before the performance improvement stagnates, suggesting that the optimal number of clusters lies in that range.
6.3. Sensitivity Analysis: Cluster Granularity and Labeling Trade-offs
We also investigate how the choice of kk—the number of clusters—impacts our model’s ability to capture and distinguish fine-grained activities. Similar to previous section, we use CASAS labels and apply majority voting per cluster to compute F1 scores. For this analysis, we report the Macro-F1 score to ensure equal weighting across all classes. This allows us to assess clustering sensitivity across the entire label distribution, preventing results from being dominated by high-frequency classes and better highlighting the model’s capacity to capture less frequent but clinically relevant activities.
Fig. 5 shows the Macro F1 scores with bootstrapped 95% confidence intervals (CIs) for Milan, Aruba, and Cairo, when $k$ varies from 10 to 100. The general trend in F1 scores indicates that utilizing more clusters generally leads to better classification performance for all three datasets, although the improvements appear to plateau after $k=30$ for Milan, and $k=20$ for Aruba and Cairo. This leads to a conclusion that increasing the number of clusters beyond 20–30 is not particularly useful. Fig. 10 in the Appendix shows what these various clusters look like on 2D tSNE projections.
@@ 7. Discussion
In this section, we discuss the real-world potential and practical deployability of DISCOVER by demonstrating its capacity to transform raw, unlabeled sensor streams into actionable behavioral insights. We first evaluate its ability to extract and monitor data-grounded longitudinal patterns (Section 7.1), a critical requirement for tracking long-term health trends in unscripted environments. We then highlight the method’s utility as an automated annotation engine (Section 7.2), which significantly reduces the human effort typically required to bootstrap new smart home deployments. Furthermore, we discuss how DISCOVER can re-annotate existing benchmark datasets at finer, more nuanced granularity levels than their native labels currently allow (Section 7.3), unlocking deeper layers of behavioral analysis. Finally, we outline current limitations and propose future research directions to further enhance the system’s robustness and adaptive capabilities (Section 7.4).
7.1. Longitudinal Analysis of Behavioral Trends
Figure 6. Longitudinal behavioral trend analysis on the CASAS Milan dataset. The comparison spans two one-week periods (Nov 25 vs. Dec 23) and highlight shift in routines. Subplots (a) and (b) display macro-shifts in native CASAS annotations, where cooking and relaxing decreased by 4-5 percentage points ($p$), while eating, sleeping and bathing increased by 2-3 pp. Subplots (c) and (d) demonstrate that our discovered clusters not only recover these high-level trends, such as increased master bedroom and bathroom activity, but provide a more nuanced, data-grounded view. For instance, DISCOVER reveals that the $3.8$ decrease in sitting in living room armchairs (CL0) is partially replaced by sitting in the TV room armchair (CL12), uncovering detailed changes in household routines through an entirely label-free approach.
The motivation for human activity recognition in smart homes extends beyond the instant detection of actions. While traditional HAR systems focus on immediate states, such as sitting, walking, or running, a critical long-term application lies in the longitudinal tracking of behavioral change. For example, from a clinical perspective, identifying shifts in daily routines is a cornerstone of health intervention research (Dunton et al., 2021; Mishra, 2019). Longitudinal tracking can validate the success of interventions aimed at reducing sedentary behavior, where success is measured by persistent, month-over-month shifts in activity distribution rather than a single day’s performance (Jekel et al., 2015). Furthermore, in the field of Mild Cognitive Impairment (MCI) and dementia research, observing gradual but persistent changes in habits, such as changes in sleep cycles or increased frequency of nocturnal bathroom visits, can serve as an early indicator of cognitive decline (Shahid et al., 2024; Attaullah et al., 2025). This makes the ability to identify and compare behavioral patterns across different periods a significant contribution to elderly care.
DISCOVER facilitates longitudinal analysis in these domains by offering a data-grounded perspective to track behavioral shifts. To demonstrate this, we compared the activity distributions of the CASAS Milan dataset across two distinct periods: one week in November and one week in December. [^1] The choice of these specific dates is dictated by data availability for the Milan dataset; we acknowledge that comparing two weeks during the holiday season is suboptimal for clinical interpretation; however, this exercise serves to validate DISCOVER’s ability to quantify longitudinal variance regardless of the specific seasonal or environmental drivers. By plotting the distribution of our 20 discovered clusters against the original CASAS annotations, we can visualize how a resident’s behavioral fingerprint evolves. Fig. 6 illustrates this phenomenon: examining the original CASAS annotations (subplots (a) and (b)), we detected several shifts in how people spent their time, measured by the percentage point (pp) change in the relative number of data samples. We see a decrease in “Cooking” ($5.3\%$), “Relaxing” ($4.8\%$), and “Working” ($1.8\%$), and an increase in “Eating” ($2.4\%$), “Sleeping” ($1.5\%$), “Bathing” ($1.7\%$), and the catch-all “Other” ($6.1\%$) categories. DISCOVER recovers similar high-level trends without requiring any ground-truth labels: subplots (c) and (d) clearly show a decrease in some kitchen movements and desk activity and an increase in master bedroom and bathroom activations, but does so without requiring any labels or supervision. More importantly, DISCOVER offers a more nuanced view specific to this household: our pipeline reveals, for example, that sitting in the living room chairs was largely replaced by time in a TV room armchair, and that kitchen shifts were characterized specifically by decreased stove use but increased refrigerator access. Furthermore, DISCOVER permits a more nuanced inspection of the non-descriptive “Other” category, which exhibited the most significant longitudinal shift in Fig. 6 with a $6.1\%$ increase. While the original CASAS annotations offer no insight into the nature of this change, DISCOVER’s comprehensive, label-free mapping captures the entirety of these sensor events, grounding previously obscure shifts in precise, spatially-anchored patterns.
In contrast to the extensively annotated CASAS dataset, the vast majority of real-world smart home environments remain unlabeled. While simpler statistical or rule-based methods may suffice for datasets like Milan, which features a sparse deployment of only 30 sensors, these approaches fail to scale as environments grow in complexity. As we introduce denser instrumentation, including contact sensors for cabinets, smart pill dispensers, and energy-load sensors, the ad-hoc configuration required to manually adjust rule-based pipelines for every unique household becomes prohibitive. DISCOVER overcomes this bottleneck by automatically extracting patterns directly from the available data, bypassing the need for standardized activity definitions. This capability is particularly vital for longitudinal tracking, where the focus shifts toward modeling within-person variability. By comparing a resident’s “old self” to their “new self,” DISCOVER can detect subtle, practically, and, hopefully, clinically significant behavioral evolutions that are unique to the individual, providing a scalable pathway toward correlating automated sensing with long-term health outcomes.
7.2. Annotation of Massive Unlabeled Ambient Sensor Datasets
@@
Smart homes have been studied extensively across academia, industry and medical fields. While fully annotated datasets such as CASAS (Cook et al., 2009), Marble (Arrotta et al., 2022), and Orange4Home (Cumin et al., 2017) are widely used in AI research, they are barely representative of the predominantly unlabeled and noisy data streams typical of most practical smart home environments. Large-scale medical initiatives such as CART (Beattie et al., 2020) (272 homes over 3 years) and TIHM (Palermo et al., 2023) (56 homes over 50 days) produce orders of magnitude more data, but remain less commonly used due to their lack of activity labels resulting from prohibitively high costs of manual annotation. With DISCOVER, it becomes feasible to automatically annotate such massive datasets. Users can choose their desired level of granularity by adjusting the number of discovered clusters, and exercise control over the annotation process via an interactive visualization tool that displays sensor activation sequences for each cluster. While such labels may not constitute true ground truth, which is inaccessible without camera data or similar information, our analysis has shown that processing with DISCOVER results in semantically meaningful activity clusters that can be labeled with high degree of agreement, thus providing insight into common everyday activities. We believe that DISCOVER can expand access to previously unlabeled data, creating opportunities for advancement in the areas such as in-home health monitoring (Morita et al., 2023), early detection of Alzheimer’s (Alberdi et al., 2018), and intelligent support for individuals with dementia (Demir et al., 2017).
7.3. Re-annotation of Existing Datasets with Finer Granularity
@@
As we have seen in the case of CASAS (Cook et al., 2012), existing datasets in the HAR domain often provide only coarse labels that capture broad activities, obscuring the rich variability of human behavior. DISCOVER addresses this shortcoming by enabling the discovery of patterns of daily living and variants within the data. For instance, while a standard dataset might label hours of activity simply as “Relaxing,” DISCOVER allows us to differentiate between distinct patterns of daily living that are otherwise collapsed into a single category. As shown in Fig. 4, the activity of “Relaxing” in this specific household is manifested through two distinct behaviors: sitting in the living room armchair versus sitting in the TV room armchair. While the specific nature of the activity remains unknown, our method grounds these annotations in the physical layout of the home, transforming a vague semantic label into a precise, spatially-anchored markers. Similarly, in meal preparation, instead of labeling the entire process as Cooking, our approach can reveal detailed steps such as time spent at the stove, visits to the pantry, and travel to/from the dining room. By re-annotating existing datasets with these finer-grained labels, DISCOVER provides deeper insights into behavioral patterns, which can be particularly valuable in clinically relevant applications and personalized intervention strategies.